Find the value of πœ•z/πœ•x and πœ•z/πœ•y at the point (1,1,1) for the surface xy +z3xβˆ’2yz = 0 by differentiating implicitly.


1
Expert's answer
2020-10-11T18:45:35-0400

z = xy +z3βˆ’2yz = 0


"\partial z/ \partial x = y+ z^3"


at point (1,1,1)

"\partial z/ \partial x" = 1 + 13 = 2


"\partial z/ \partial y = x +2z"


at point 1,1,1


"\partial z/ \partial y" = 1 + 2*1 = 3